OpenStudy (anonymous):

find the exact value of tan^2( csc^-1(7/4))

OpenStudy (anonymous):

Any idea??

OpenStudy (anonymous):

Hint: \(csc \theta =\dfrac{1}{sin \theta}\\sin^-({somenumber})=\theta\)

OpenStudy (anonymous):

so...

OpenStudy (anonymous):

apply the hint, that's it